# Hyperbolic regularization effects for degenerate elliptic equations

### v1

@@ 7.2. Characterization of blow-up limits
Assume (1.9) and (7.3). Let $w\in\operatorname{Lip}(B_{1},\mathbb{R}^{2})$ solve $Dw\in K$ a.e., and $x_{0}\in B_{1}$. Any blow-up limit $w_{\infty}\in\mathcal{B}(w)(x_{0})$ can only be singular at 0 and $Dw_{\infty}$ is 0-homogenenous.
The proof of Proposition 7.2 uses the structure of characteristic lines established in Corollary 6.6, combined with Theorem C and the following lemma.
@@
Since $JDu$ is divergence-free, this identity is preserved after substracting a linear function from $u$, so we may without loss of generality assume that $u=0$ on $B_{1}^{-}$, that is, $f=0$ on $\pi/2,3\pi/2]$. Choosing radial test functions $\varphi(re^{i\theta})=\psi(r)$ with $\psi\in C_{c}^{2}([0,1))$ such that $\psi^{\prime}(0)=\psi^{\prime\prime}(0)=0$ and $\psi(0)=1$, (7.6) becomes
@@
As argued above, by Theorem 7.1 and Theorem 1.3, under assumptions (1.9)-(7.3), the blow-up limit $w_{\infty}$ satisfies $Dw_{\infty}\in\Gamma$ a.e. in $\mathbb{R}^{2}$, where $\Gamma=\gamma(I)$ is a compact connected $C^{1}$ curve satisfying (1.9), as in § 6. We can assume that $0\in\operatorname{Sing}(w_{\infty})$, as otherwise $x_{0}\in\operatorname{Reg}(w)$ and $w_{\infty}$ is affine by Theorem A. Letting $\theta=\gamma^{-1}(Dw_{\infty})$ and $\Psi$ as in Lemma 6.4, we have therefore that the field of characteristic lines $v=\Psi\circ\theta$ must be of the form (6.20). On any subset $A\subset\mathbb{R}^{2}\setminus\operatorname{Sing}(w_{\infty})$ on which $v$ is constant, say $v=v_{0}$, $Dw_{\infty}$ must belong to $\Psi^{-1}(\{v_{0}\})$, a totally disconnected set by Lemma 6.4, and therefore be constant. Combining this with the structure (6.20) of the map $v$, we deduce that $Dw_{\infty}$ is either constant, or 0-homogeneous, or that it has two discontinuity points, 0 and $y_{0}\neq 0$.
To prove Proposition 7.2 we just need to rule out the latter case. To do so, by Theorem A it suffices to show that there exists a linear blow-up of $w_{\infty}$ at $y_{0}$. Let $W$ be any such blow-up. Due to the structure of the characteristic lines of $w_{\infty}$ (6.20), we infer that $DW$ is 0-homogeneous and constant in $P_{\alpha}\doteq\{(x,e^{i\alpha})<0\}$ for some $e^{i\alpha}\in\mathbb{S}^{1}$. Up to a rotation, we can assume without loss of generality that $\alpha=0$. Hence its first component, $u\in C^{1}(\mathbb{R}^{2}\setminus\{0\})$, is $1$-homogeneous and coincides with a linear function on $P_{0}$. In addition, since $w_{\infty}\in\mathcal{B}(w)(x_{0})$, then $y_{0}\notin\operatorname{spt}(\mathcal{D}(w_{\infty}))$ by Corollary 5, and hence for the same reason $\mathcal{D}(u,u)=0$ in $\mathbb{R}^{2}$. Combining this information with the structure of $u$ we can infer from Lemma 7.2 that $u$ is linear, and hence $y_{0}\notin\operatorname{Sing}(w_{\infty})$, a contradiction. ∎
@@
Let $x_{0}\in\operatorname{Sing}(w)$ such that $B_{\delta}(x_{0})\subset\Omega$, and assume that $x_{0}$ is not isolated. Then there exists a sequence $x_{j}\in\operatorname{Sing}(w)\setminus\{x_{0}\}$ such that $x_{j}\to x_{0}$. Let $r_{j}=|x_{j}-x_{0}|\to 0$ and consider the maps $w_{j}\doteq w_{r_{j},x_{0}}$, see (1.13). They satisfy $0,y_{j}\in\operatorname{Sing}(w_{j})$ where $y_{j}=(x_{j}-x_{0})/r_{j}\in\mathbb{S}^{1}$. Consider a (non relabeled) subsequence such that $w_{j}\to w_{\infty}\in\mathcal{B}(w)(x_{0})$ and $y_{j}\to y_{\infty}\in\mathbb{S}^{1}$. By Theorem A and the assumption $\varepsilon_{*}(K)>0$ we have $0,y_{\infty}\in\operatorname{Sing}(w_{\infty})$, in contradiction with Proposition 7.2. ∎
Combining Corollary 7.2 with Propositions 7.1 and 6.4, we deduce the following partial regularity result.
@@
Finally, recalling that a compact connected $C^{1}$ curve $\Gamma\subset\mathbb{R}^{2\times 2}$ satisfying (1.9) has $\varepsilon_{*}(\Gamma)=+\infty$ if it is simply connected or if its first-row projection is the boundary of a strictly convex open set, we see that Theorem E follows from Corollary 7.2 and Proposition 2.2.

### v2

@@ 9.2. Characterization of blowup limits
Assume (2.4) and (9.5). Let $w\in\operatorname{Lip}(B_{1},\mathbb{R}^{2})$ solve $Dw\in K$ a.e., and $x_{0}\in B_{1}$. Any blowup limit $w_{\infty}\in\mathcal{B}(w)(x_{0})$ can only be singular at 0 and $Dw_{\infty}$ is 0-homogenenous.
We start by ruling out specific singularities if the weak Hessian determinant vanishes.
@@
Since $JDu$ is divergence-free, this identity is preserved after subtracting a linear function from $u$, so we may without loss of generality assume that $u=0$ on $B_{1}^{-}$, that is, $f=0$ on $\pi/2,3\pi/2]$. Choosing radial test functions $\varphi(re^{i\theta})=\psi(r)$ with $\psi\in C_{c}^{2}([0,1))$ such that $\psi^{\prime}(0)=\psi^{\prime\prime}(0)=0$ and $\psi(0)=1$, (9.8) becomes
@@
As argued above, by Theorem 9.2 and Theorem 2.4, under assumptions (2.4)-(9.5), the blowup limit $w_{\infty}$ is either linear or it satisfies $Dw_{\infty}\in\Gamma$ a.e. in $\mathbb{R}^{2}$, where $\Gamma=\gamma(I)$ is a compact connected $C^{1}$ curve satisfying (2.4), as in § 8. We can assume we are in the latter case, as otherwise there is nothing to show, and for the same reason we can assume that $0\in\operatorname{Sing}(w_{\infty})$, since otherwise $x_{0}\in\operatorname{Reg}(w)$ and $w_{\infty}$ is affine by Theorem 6.2. Letting $\theta=\gamma^{-1}(Dw_{\infty})$ and $\Psi$ as in Lemma 8.4, we have therefore that the field of characteristic lines $v=\Psi\circ\theta$ must be of the form (8.20). On any subset $A\subset\mathbb{R}^{2}\setminus\operatorname{Sing}(w_{\infty})$ on which $v$ is constant, say $v=v_{0}$, $Dw_{\infty}$ must belong to $\Psi^{-1}(\{v_{0}\})$, a totally disconnected set by Lemma 8.4, and therefore be constant. Combining this with the structure (8.20) of the map $v$, we deduce that $Dw_{\infty}$ is either constant, or 0-homogeneous, or that it has two discontinuity points, 0 and $y_{0}\neq 0$.
To prove Proposition 9.2 we just need to rule out the latter case. To do so, by Theorem 6.2 it suffices to show that there exists a linear blowup of $w_{\infty}$ at $y_{0}$. Let $W$ be any such blowup. Due to the structure of the characteristic lines of $w_{\infty}$ (8.20), we infer that $DW$ is 0-homogeneous and constant in $P_{\alpha}\doteq\{(x,e^{i\alpha})<0\}$ for some $e^{i\alpha}\in\mathbb{S}^{1}$. Up to a rotation, we can assume without loss of generality that $\alpha=0$. Hence its first component, $u\in C^{1}(\mathbb{R}^{2}\setminus\{0\})$, is $1$-homogeneous and coincides with a linear function on $P_{0}$. In addition, since $w_{\infty}\in\mathcal{B}(w)(x_{0})$, then $y_{0}\notin\operatorname{spt}(\mathcal{D}(w_{\infty}))$ by Corollary 7, and hence for the same reason $\mathcal{D}(u,u)=0$ in $\mathbb{R}^{2}$. Combining this information with the structure of $u$ we can infer from Lemma 9.2 that $u$ is linear, and hence $y_{0}\notin\operatorname{Sing}(w_{\infty})$, a contradiction. ∎
@@
Let $x_{0}\in\operatorname{Sing}(w)$ such that $B_{\delta}(x_{0})\subset\Omega$, and assume that $x_{0}$ is not isolated. Then there exists a sequence $x_{j}\in\operatorname{Sing}(w)\setminus\{x_{0}\}$ such that $x_{j}\to x_{0}$. Let $r_{j}=|x_{j}-x_{0}|\to 0$ and consider the maps $w_{j}\doteq w_{r_{j},x_{0}}$, see (6.1). They satisfy $0,y_{j}\in\operatorname{Sing}(w_{j})$ where $y_{j}=(x_{j}-x_{0})/r_{j}\in\mathbb{S}^{1}$. Consider a (non relabeled) subsequence such that $w_{j}\to w_{\infty}\in\mathcal{B}(w)(x_{0})$ and $y_{j}\to y_{\infty}\in\mathbb{S}^{1}$. By Theorem 4.1 and the assumption $\varepsilon_{*}(K)>0$ we have $0,y_{\infty}\in\operatorname{Sing}(w_{\infty})$, in contradiction with Proposition 9.2. ∎
Combining Corollary 9.2 with Propositions 9.1 and 8.4, we deduce the following regularity result:
@@
Finally, recalling that $\Gamma\subset\mathbb{R}^{2\times 2}$ has $\varepsilon_{*}(\Gamma)=+\infty$ if its first-row projection is the boundary of a strictly convex open set, or if it is a compact and simply connected $C^{1}$ curve $\Gamma\subset\mathbb{R}^{2\times 2}$ satisfying (2.4), we see that Theorem 9.1 follows from Corollary 9.2 and Proposition 3.2.